OpenStudy (anonymous):

how would i find y for (8y+4)+(4y+8)=15y-9 @Compassionate

OpenStudy (anonymous):

i forgot what PEMDAS is

OpenStudy (compassionate):

(8y+4)+(4y+8)=15y-9 PEMDAS: Peranthesis, Exponents, Multiplication, Division, Addition, Subtraction. It's called the Order of Operations. You want to simplify Parenthesis, then exponents, then multiplication, and what not. You move from left to right. Okay, so first you want to solve inside the parenthesis. You see, you cannot simplify it anymore, so we want to just ignore them. 8y + 4 + 4y + 8 = 15y - 9 Okay, so there is no parenthesis, there are no exponents, multiplication, OR division, so we just add like terms. 12y + 12. This is because we added the like terms 8y and 4y, and 8 + 4. So now we have: 12y + 12 = 15y - 9 Now the equation is asking us to solve for y, and remember, we solve from the left to the right, so we just perform the opposite action now and solve it like any other normal linear equation. Subtract 12y 12y - 12y + 12 = 15y - 12y - 9 12 = 3y - 9 Now add 9. 21 = 3y Now, do the opposite, divide 3. 21/3 = y y 7 = y Or: y = 7 . Does that make sense, hun?

OpenStudy (compassionate):

Also, \(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\)

OpenStudy (anonymous):

Yes it makes perfect sense thank you so much and thank you for the welcome! :D

OpenStudy (compassionate):

If you need absolutely anything else, feel free to tag me, or anyone else you find helpful such as Ambassadors and Moderators.

OpenStudy (anonymous):

Thank you I will certainly